Royal Lawncare Company produces and sells two packaged products—Weedban and Greengrow. Revenue and cost information relating to the products follow:

Product

Weedban Greengrow
Selling price per unit $ 8.00 $ 38.00
Variable expenses per unit $ 2.40 $ 11.00
Traceable fixed expenses per year $ 137,000 $ 48,000

Common fixed expenses in the company total $102,000 annually. Last year the company produced and sold 37,500 units of Weedban and 16,000 units of Greengrow.

Required:

Prepare a contribution format income statement segmented by product lines.
Answer #1
Total Weedban Green grow
Sales 908,000 (37500*8) = 300,000 (16000*38) = 608,000
Variable cost 266,000 (37500*2.40)= 90,000 (16000*11) = 176,000
Contribution margin 642,000 210,000 432,000
Traceable fixed expenses 185,000 137,000 48,000
Segment margin 457,000 73,000 384,000
Common fixed costs 102,000
Net operating income 355,000